Use this problem for #3 and 4.
Lucy rides her bike 4 km/h faster than Kim.
Lucy travels east for 2 hours.
Kim travels west for 2 hours.
They are then 80 km apart.
How many km/h is Kim traveling?

Answer by fractalier(6550)

Remember D = RT, distance = rate*time.
Let Kim's rate be r.
Then Lucy's rate is r + 4.
They both travel for 2 hours, so t = 2.
Their total distance is 80 km, so we have
2r + 2(r + 4) = 80
2r + 2r + 8 = 80
4r + 8 = 80
4r = 72
r = 18 km/h, Kim's rate.